OpenStudy (kimo_indie):

You want to prepare a test that contains 20 questions worth 50 points overall. How many 2-point and 4-point questions can you include in the test? a) 5, 10 b) 10, 10 c) 12, 8 d) 15, 5

OpenStudy (bibby):

treat them as ordered pairs. you want 2x+4y=50 x+y=20 so we can cancel the choices where there aren't 20 questions in total (a: 5+10=15) now find which one of b,c, or d fits 2x+4y=50

OpenStudy (kimo_indie):

so d then?

OpenStudy (kimo_indie):

thnx

OpenStudy (bibby):

yeah